OpenStudy (anonymous):

equation in slope- intercept, point slope, or standard form for thr line with the given information please explain why you chose form. 1. passes through (-1,4) and (-5,2)

OpenStudy (anonymous):

slope= (2-4)/(-5-(-1))=-2/-4=1/2